Hi
20% of the toys in a store are dolls. There are 286 more cars than dolls. The remaining 650 items are balls. How many toys are in the store.
Thanks

Answer by ikleyn(52756)

.

Your equation is

    0.2x + (0.2x + 286) + 650 = x.


Solve for x

    286 + 650 = x - 0.2x - 0.2x

       936    =     0.6x

        x     =     936/0.6 = 1560.


ANSWER.  There are 1560 toys in the store.

Solved.